On January 1, 2021, HGC Camera Store adopted the dollar-value LIFO retail inventory method. Inventory transactions at both cost and retail, and cost indexes for 2021 and 2022 are as follows:

2021 2022
Cost Retail Cost Retail
Beginning inventory $ 39,200 $ 56,000
Net purchases 92,600 116,000 $ 104,166 $ 128,400
Freight-in 2,800 3,300
Net markups 14,000 9,600
Net markdowns 2,800 3,000
Net sales to customers 113,355 117,120
Sales to employees (net of 20% discount) 3,300 5,280
Price Index:
January 1, 2021 1.00
December 31, 2021 1.06
December 31, 2022 1.10


Required:
Estimate the 2021 and 2022 ending inventory and cost of goods sold using the dollar-value LIFO retail inventory method.
Answer #1


For 2022: 5280 / 0.80 = 6600 - 5280 = 1320 = Employee Discounts Beginning inventory Plus: Net Purchases Cost 43970 104166 330